### v1

# Low-Latency and Low-Complexity MLSE for Short-Reach Optical Interconnects

Abstract
To meet the high-speed, low-latency, and low-complexity demand for optical interconnects, simplified layered 2-step maximum likelihood sequence estimation (L2S-MLSE) is proposed in this paper. Simplified L2S-MLSE combines computational simplification and reduced state in L2S-MLSE. L2S-MLSE with a parallel sliding block architecture reduces latency from linear order to logarithmic order. Computational simplification reduces the number of multipliers from exponential order to linear order. Incorporating the reduced state with computational simplification further decreases the number of adders and comparators. The simplified L2S-MLSE is evaluated in a 112-Gbit/s PAM4 transmission over 2-km standard single-mode fiber. Experimental results show that the simplified L2S-MLSE significantly outperforms the FFE-only case in bit error ratio (BER) performance. Compared with simplified 1-step MLSE, the latency of simplified L2S-MLSE is reduced from 34 delay units in linear order to 7 delay units in logarithmic order. The simplified scheme in L2S-MLSE reduces the number of variable multipliers from 512 in exponential order to 33 in linear order without BER performance deterioration, while reducing the number of adders and comparators to 37.2% and 8.4%, respectively, with nearly identical BER performance.
Index Terms:
Maximum likelihood sequence estimation, low latency, low complexity, optical interconnects.
I Introduction
[…]
In the DD-FTN scheme, MLSE not only eliminates the residual ISI from FFE, but also eliminates the ISI introduced by the post-filter. However, the performance improvement achieved by MLSE comes with high computational complexity of the Viterbi algorithm. The research about reducing the computational complexity of MLSE is mainly divided into two categories: reducing the number of states and reducing the number of multipliers. To reduce the number of states, the output of FFE can be used as a pre-decision value to reduce the number of candidate states in MLSE [hoon, jing, Hiroki]. Through the pre-decision value, the candidate states with high probability are reserved, while the other states are discarded to reduce the trellis size. Moreover, the decision region whether the reduced-state MLSE is applied can be iteratively determined by error rate [li]. To reduce the number of multipliers, in the branch metric (BM) calculation of MLSE, the lookup table is used to replace the convolution calculation [dai]. The combination of lookup table and reduced state can further reduce the computational complexity [jing2, li2]. The piecewise linear formula or absolute value can also be used to replace the squaring operation in BM calculation [dai, cai].
Besides the complexity, during the add-compare-select (ACS) calculation of MLSE, the current BM should continuously add the previous accumulated metrics in series. This serial calculation structure leads to the continuous increase of latency. With the rapid development of latency-constrained services, such as augmented reality, virtual reality, autonomous driving, and real-time gaming, the demand for low-latency data center optical interconnects is becoming urgent [latency1, latency2, latency3]. Fettweis and Meyr first proposed the look-ahead M-step technique, which can effectively break the latency bottleneck of ACS operation in the conventional Viterbi algorithm [M1, M2, M3]. Inspired by the look-ahead M-step technique, the layered look-ahead Viterbi decoding architecture is used to reduce the latency [M4, M5]. Although the look-ahead M-step technique efficiently decreases the latency, its computational complexity still remains difficult to accept, especially when the modulation order of the signal, the number of channel coefficients in the post-filter, and the number of steps in the M-step are increased.
To meet the high-speed, low-complexity, and low-latency demand for optical interconnects, the simplified layered 2-step maximum likelihood sequence estimation (L2S-MLSE) is proposed in this paper. The latency advantage of the layered 2-step structure and the complexity reduction of simplified L2S-MLSE are analyzed in detail. The bit error ratio (BER) performance of simplified L2S-MLSE is experimentally verified in a 112-Gb/s PAM4 transmission over 2-km SSMF. The main contributions of this paper are as follows:
  • The proposed simplified L2S-MLSE combines the computational simplification and the reduced state in L2S-MLSE. The latency is reduced from linear to logarithmic order through L2S-MLSE with a parallel sliding block architecture. Computational simplification reduces the number of multipliers from exponential to linear order. Incorporating the reduced state further decreases the number of adders and comparators.
  • Experimental results show that compared with simplified 1-step MLSE (1S-MLSE), the latency of simplified L2S-MLSE is decreased from 34 delay units in linear order to 7 delay units in logarithmic order without performance deterioration. Compared with L2S-MLSE, the simplified L2S-MLSE reduces the number of variable multipliers from 512 in exponential order to 33 in linear order, while the number of adders and comparators is reduced to 37.2% and 8.4%, respectively.
The rest of this paper is organized as follows. The principle of parallel 1S-MLSE is introduced in Section II. The parallel L2S-MLSE with low-latency property is described in Section III. In Section IV, the proposed simplified L2S-MLSE is described. In Section V, we present the experimental setup of 112-Gb/s PAM4 transmission over 2-km SSMF. In Section VI, we analyze the parameter settings, BER performance, latency and computational complexity of simplified L2S-MLSE in the experiment. Finally, the paper is concluded in Section VII.
Figure 1: (a) Block diagram of three steps in MLSE. (b) Trellis diagram of PAM4 signal with 2-tap channel response. Sel.: Select.
$PM_{n}(s_{n})=\min\{$ $PM_{n-1}(-3)+BM_{n}(-3,s_{n}),PM_{n-1}(-1)+BM_{n}(-1,s_{n}),$ (3)
$PM_{n-1}(1)\hskip 8.0pt+BM_{n}(1,s_{n}),\hskip 8.0ptPM_{n-1}(3)\hskip 8.0pt+BM_{n}(3,s_{n})\}$
II Parallel 1S-MLSE
II-A Signal Processing Procedure in MLSE
MLSE employs the Viterbi algorithm to search for the most-likely state transition sequence in the state trellis [viterbi1, viterbi2], which is usually placed after the FFE and post-filter [zhong]. The FFE can compensate for ISI at the cost of noise enhancement. The enhanced noise and the equalized signal are suppressed by the post-filter. Meanwhile, the post-filter can approximate the original channel response with the known shortened coefficients. Then, the signal filtered by the post-filter is fed into MLSE with the known ISI coefficients. The signal processing procedure of MLSE includes three steps, as shown in Fig. 1(a). The three steps are BM calculation, ACS, and survivor path selection.
Figure 2: Sliding block architecture for (a) serial MLSE and (b) parallel MLSE. R: Received symbols; O: Overlapped symbols; D: Data symbols.
1) BM Calculation: In the BM calculation, the BM at time index $n$ can be calculated as
$BM_{n}=\left|y_{n}-\sum_{k=0}^{L-1}{\alpha_{k}s_{n-k}}\right|^{2}$ (1)
where $s_{n}$ and $y_{n}$ are the transmitted signal and received signal at time index $n$. The vector $\bm{\alpha}$ represents the coefficients of the channel response, which includes the elements from $\alpha_{0}$ to $\alpha_{L-1}$, where $\alpha_{0}$ is equal to 1. For the PAM-$M$ signal, each transmitted signal $x_{n}$ at time index $n$ has $M$ possible different states. With $L$ coefficients of the channel response, the trellis has $M^{L-1}$ different states at each time index, and each previous state at time index n1n-1 extends $M$ different branches to the current state at time index $n$. Thus, the total number of BMs from time index n1n-1 to time index $n$ is $M^{L}$. To reduce computational complexity, a 2-tap post-filter ($L=2$) is often used, as it minimizes the number of BMs. For PAM4 modulation with a 2-tap channel response, the BM is calculated as
$BM_{n}(s_{n-1},s_{n})=\left|y_{n}-\left(\alpha s_{n-1}+s_{n}\right)\right|^{2}$ (2)
where $s_{n-1}\in\{-3,-1,1,3\}$ and $s_{n}\in\{-3,-1,1,3\}$. $BM_{n}(s_{n-1},s_{n})$ represents the BM from state $s_{n-1}$ to state $s_{n}$ at time index $n$. Eq. (2) totally includes 16 variable multipliers, 1 constant multiplier for $3\times\alpha$, and 32 adders.
2) ACS Operation: The objective of ACS is to find the best path at each state. Firstly, the accumulated value of previous BMs, which is named as path metric (PM) for each state, is added with the current BM extended from the corresponding state. Secondly, as shown in Fig. 1(b), four different paths converge to the same state, and the path with the minimal PM is chosen at each state (an example is illustrated as a solid line). In conclusion, the ACS computing process can be expressed as Eq. (3), where $PM_{n}(s_{n})$ represents the PM of state $s_{n}$ at time index $n$. The minimal PM value at each state serves as the previous accumulated value for the next ACS calculation. Each ACS unit, as shown in Eq. (3), includes 4 adders and 3 comparators. In this 4-to-1 selection, the first two comparators are used for two 2-to-1 comparisons. Their results are input into the third comparator to obtain the 4-to-1 result. The signal $s_{n}$ includes 4 states, thus the total 4 ACS units at time index $n$ include 16 adders and 12 comparators.
3) Survivor Path Selection: When the last 4 ACS calculations are finished, 4 different paths are reserved. Therefore, in the survivor path selection, the path with the minimal PM is chosen from the last 4 paths. This 4-to-1 selection still requires 3 comparators. Finally, the output of MLSE is determined based on the path with the minimal PM.
The received symbols calculated in the successive ACS should have a finite length. In the conventional MLSE, when the receiver receives a frame of symbols, the received symbols are processed by serially segmented sliding blocks. As illustrated in Fig. 2(a), each overlapped sliding block of $N$ symbols includes $R$ received symbols followed by $O$ overlapped symbols [TB]. In each sliding block, when the ACS calculation completes at the last overlapped symbol, the final survivor path with $N=R+O$ symbols is obtained. Then, the $R$ received symbols are decoded into $D$ data symbols. Only the results of data symbols are reserved, and the overlapped symbols are removed. The PM of $D$ data symbols is transmitted to the next sliding block as the initial accumulated BM.
II-B Parallel 1S-MLSE: Analysis of Latency and Complexity
In practical MLSE implementation, the received symbols should be processed in parallel [sliding]. As shown in Fig. 2(b), the decision of the state at time index $n$ only depends on the trellis from nOn-O to $n+O$ [sliding]. With the overlap before and behind the data, the received symbols in different blocks can be processed in parallel. Therefore, the received symbols in different blocks are processed through parallel BM calculation, parallel ACS, and parallel survivor path selection. In our paper, all the MLSE algorithms described in the following are based on the parallel sliding block architecture. For the parallel MLSE described in this section, although the received symbols in different blocks are processed in parallel, the symbols within the same block are still processed one after another serially in ACS. To distinguish from the 2-step processing method described in Section III, the parallel MLSE described in this section is termed 1S-MLSE, as symbols within the same block are still processed one after another by 1-step.
The architecture of 1S-MLSE is shown in Fig. 3. One of the parallel sliding blocks with the symbol length of $N=O+R+O$ is analyzed. The BM of all the $N$ received symbols can be calculated in parallel. The PM of the previous time must be transferred to the next time during the ACS operation (the initial PM for the first ACS is zero). Consequently, the prepared BM must wait sequentially as illustrated in Fig. 3. For the block with $N$ received symbols, the latency of ACS should be $N$ delay units, which increases with the increase of block length. Both the BM calculation and survivor path selection introduce the latency of $1$ delay unit, resulting in the overall latency of $N+2$ delay units for 1S-MLSE.
Finally, the computational complexity of 1S-MLSE with the symbol length of $N$ is studied. For the BM calculation, the channel response coefficients should be configurable. Considering that the channel response changes slowly, the equation $\alpha s_{n-1}+s_{n}$ can be calculated only once for every $N$ symbols. Therefore, the computational complexity of BM calculation includes $16N$ variable multipliers, 1 constant multiplier, and $16N+16$ adders. For the ACS calculation, 16(N1)16(N-1) adders and $12N$ comparators are included. In the survivor path selection, the path with the minimal PM is chosen from the last 4 paths, which requires 3 comparators. In conclusion, the overall computational complexity of 1S-MLSE includes $16N$ variable multipliers, 1 constant multiplier, $32N$ adders, and $12N+3$ comparators.
Figure 3: Signal processing architecture for 1S-MLSE in an NN-symbol sliding block based on serial ACS operation.
III Low-Latency L2S-MLSE
In the 1S-MLSE, symbols in the ACS calculation should be processed one after another. The latency of ACS for $N$ received symbols in a sliding block is $N$ delay units, which is increased with the increase of the value $N$. Inspired by the look-ahead $M$-step technique [M1, M2, M3, M4, M5], L2S-MLSE with parallel sliding block architecture can be applied to reduce the latency. Fig. 4(a) illustrates the processing details of L2S-MLSE with $N$ received symbols. The main difference between 1S-MLSE and L2S-MLSE is in the process of ACS. In the 1S-MLSE, one group of $16$ BMs from time index n1n-1 to time index $n$ is processed by 4 ACS units at a time. In the L2S-MLSE, two successive steps of $32$ BMs from time index n1n-1 to time index $n+1$ are processed by 16 ACS units in the first layer. After the first layer, the output of two groups of accumulated BM is still processed by 16 ACS units in the next layer. Finally, the last two groups of accumulated BM are input into 16 ACS units, and the survivor path is found from the outputs of these 16 ACS units. The L2S-MLSE trellis in Fig. 4(b) illustrates an example with a symbol length $N=8$. Owing to the parallel processing, it is obvious that the latency of ACS is reduced from $N$ delay units to $\log_{2}(N)$ delay units.
Figure 4: (a) Signal processing architecture for L2S-MLSE in an NN-symbol sliding block based on layered parallel ACS operation. BM: branch metric; ACS: add-compare-select. (b) Example of an L2S-MLSE trellis with $N=8$. L1: First layer; L2: Second layer; L3: Third layer.
$BM_{n+1}(s_{n-1},s_{n+1})=\min\{$ $BM_{n}(s_{n-1},-3)+BM_{n+1}(-3,s_{n+1}),BM_{n}(s_{n-1},-1)+BM_{n+1}(-1,s_{n+1}),$ (4)
$BM_{n}(s_{n-1},1)\hskip 8.0pt+BM_{n+1}(1,s_{n+1}),\hskip 8.0ptBM_{n}(s_{n-1},3)\hskip 7.2pt+BM_{n+1}(3,s_{n+1})\}$
1) BM Calculation and ACS of the First Layer: The signal processing procedure of L2S-MLSE still includes BM calculation, ACS, and survivor path selection. The BM calculation of L2S-MLSE is the same as that of 1S-MLSE. The details of the ACS calculation of the first layer are shown in Fig. 5. In the first layer, the BMs from time index n1n-1 to time index $n+1$ that share the same initial and final states are summed together. For the 16 ACS units that share the same initial and final states, the calculation can be expressed as Eq. (4). Since all four paths in Eq. (4) start from the same state, their BMs are comparable regardless of the previous PM. In each ACS unit, 4 adders and 3 comparators are included. In the overall 16 ACS units depicted in Fig. 5, 64 adders and 48 comparators are included. In the first layer with $N$ symbols, the overall groups of 16 ACS units are $N/2$. In conclusion, with symbol length of $N$, the overall computational complexity of the BM calculation and ACS operation in the first layer includes $16N$ variable multipliers, $1$ constant multipliers, $48N+16$ adders, and $24N$ comparators.
2) ACS of the Other Layers and Survivor Path Selection: After the first layer, the accumulated BM are transferred layer by layer. The overall groups of 16 ACS units are decreased from the first layer of $N/2$ to the following layers of $N/4$, $N/8$, and so on. In the last layer, the survivor path is found from the outputs of the last group of 16 ACS units. The survivor path selection requires 15 comparators and introduces the latency of $1$ delay unit. Therefore, the overall latency of L2S-MLSE is $\log_{2}(N)+2$ delay units. The overall computational complexity of L2S-MLSE is $16N$ variable multipliers, 1 constant multiplier, 80N4880N-48 adders, and 48N3348N-33 comparators. Compared with 1S-MLSE, the latency of L2S-MLSE is reduced from linear order to logarithmic order at the cost of increased adders and comparators.
IV Simplified L2S-MLSE
Since the latency of L2S-MLSE is reduced at the cost of complexity, simplified L2S-MLSE is proposed in this section. Simplified L2S-MLSE consists of two parts: computational simplification and reduced state. Computational simplification is proposed to reduce the number of multipliers. Reduced state is combined with computational simplification, further reducing the number of adders and comparators.
IV-A Computational Simplification
1) BM Calculation and ACS of the First Layer: To reduce the number of multipliers, the computational simplification is proposed, which combines the common terms in the BM calculation and addition operations of the first layer. The 64 expansion equations of the BM calculation and addition operations from time index n1n-1 to time index $n+1$ in the first layer are shown in Appendix A. These 64 expansion equations can be expressed as the sum of vectors $\bf{A}$}, $\bf{B}$}, $\bf{C}$}, and $\bf{D}$}. Table A.I in Appendix A provides the detailed definitions of these vectors. Both the vector ${\bf{A}}=[{A_{1}},{A_{2}},...,{A_{16}}]$ and the vector ${\bf{C}}=[{C_{1}},{C_{2}},...,{C_{16}}]$ are related to $\alpha$ and constant values. Both the vector ${\bf{B}}=[{B_{1}},{B_{2}},...,{B_{16}}]$ and the vector ${\bf{D}}=[{D_{1}},{D_{2}},...,{D_{16}}]$ are related to the received symbol $y$.
Figure 5: Illustration of 2-step trellis from time index n1n-1 to time index $n+1$ in the first layer of L2S-MLSE.
Through combining the common terms, the vectors $\bf{A}$} and $\bf{C}$} include 1 variable multiplier for $\alpha^{2}$, 4 constant multipliers for $10\alpha^{2}$, $18\alpha^{2}$, $12\alpha$, $36\alpha$, and additional 31 adders. Multiplication with powers of 2 can be achieved through bit-shift operations, thereby avoiding the use of multipliers. For the calculation of $\bf{B}$} and $\bf{D}$}, $\bf{B}$} can be decomposed into ${\bf{M}}+{\bf{N}}+{\bf{J}}$, while $\bf{D}$} can be decomposed into ${\bf{M}}+{\bf{N}}+{\bf{K}}$. $\bf{M}$} corresponds to the common term $y_{n}+\alpha y_{n+1}$ and its multiples. $\bf{N}$} corresponds to the common term $\alpha y_{n}$ and its multiples. $\bf{J}$} is equal to the common term $6y_{n+1}$ for $\bf{B}$}, and $\bf{K}$} is equal to the common term $2y_{n+1}$ for $\bf{D}$}. Then, the calculation of $\bf{B}$} and $\bf{D}$} can be represented as 5 steps:
  • Computing 2 variable multipliers: $\alpha y_{n}$ and $\alpha y_{n+1}$;
  • Computing 1 adder: common term $y_{n}+\alpha y_{n+1}$;
  • Computing 3 constant multipliers: $6y_{n}+6\alpha y_{n+1}$, $6\alpha y_{n}$ and $6y_{n+1}$;
  • Computing 4 adders for ${\bf{N}}+{\bf{J}}$ and 4 adders for ${\bf{N}}+{\bf{K}}$;
  • Computing 16 adders for ${\bf{M}}+{\bf{N}}+{\bf{J}}$ and 16 adders for ${\bf{M}}+{\bf{N}}+{\bf{K}}$.
Therefore, through combining the common terms, the vectors $\bf{B}$} and $\bf{D}$} include 2 variable multipliers, 3 constant multipliers, and 41 adders. The computation of $\bf{B}$} and $\bf{D}$} can be performed simultaneously with that of $\bf{A}$} and $\bf{C}$}. In ACS, 64 adders are required from Eq. (A.1) to Eq. (A.16) shown in Appendix A, while 48 comparators are utilized after the addition operations. In conclusion, with computational simplification, the overall complexity of BM calculation and ACS operation in the first layer includes $N+1$ variable multipliers, $3N/2+4$ constant multipliers, $105N/2+31$ adders, and $24N$ comparators.
2) ACS of the Other Layers and Survivor Path Selection: The difference between L2S-MLSE with and without computational simplification only exists in the first layer, since the second layer can only utilize the computationally simplified results from the first layer. From the second layer to the last layer, the computing process of L2S-MLSE with computational simplification is identical to that of L2S-MLSE without computational simplification, which preserves the same latency and computational complexity.
To make a fair comparison, the computational simplification is also applied to 1S-MLSE. The expansion equations of the BM calculation from time index n1n-1 to time index $n$ are shown in Appendix B. The expansion equations can be expressed as the sum of vectors $\bf{E}$} and $\bf{F}$}. Table B.I in Appendix B provides the detailed definitions of these vectors. By combining common terms, the vector ${\bf{E}}=[{E_{1}},{E_{2}},...,{E_{8}}]$ includes 1 variable multiplier for $\alpha^{2}$, 3 constant multipliers for $9\alpha^{2}$, $6\alpha$, $18\alpha$, and 12 adders. The vector ${\bf{F}}=[{F_{1}},{F_{2}},...,{F_{8}}]$ includes 1 variable multiplier for $\alpha y_{n}$, 2 constant multipliers for $6\alpha y_{n}$ and $6y_{n}$, plus 8 adders. 16 adders are needed for the addition between $\bf{E}$} and $\bf{F}$}, while 12 comparators are utilized after the addition operations. After the BM calculation with computational simplification, ACS and survivor path selection are the same as those of 1S-MLSE without computational simplification, which preserves the same latency and computational complexity.
Figure 6: BER performance of 1S-MLSE, L2S-MLSE, and their counterparts with computational simplification (CS).
In conclusion, the overall computational complexity of 1S-MLSE with computational simplification is $N+1$ variable multipliers, $2N+3$ constant multipliers, 40N440N-4 adders, and $12N+3$ comparators, while that of L2S-MLSE with computational simplification is $N+1$ variable multipliers, $3N/2+4$ constant multipliers, 169N/233169N/2-33 adders, and 48N3348N-33 comparators. Both 1S-MLSE with computational simplification and L2S-MLSE with computational simplification use the same number of variable multipliers. Because constant multipliers and comparators can be implemented by adders, variable multipliers occupy far more hardware resources than constant multipliers, adders, or comparators. When using computational simplification, the number of variable multipliers is reduced from $O(NM^{L})$ to $O(N)$. The BER performance of 1S-MLSE, 1S-MLSE with computational simplification, L2S-MLSE, and L2S-MLSE with computational simplification is shown in Fig. 6. All four algorithms have the same BER performance. By combining common items, the calculation accuracy remains unchanged, ensuring the computational simplification does not degrade the BER performance.
IV-B Simplified L2S-MLSE with the Combination of Computational Simplification and Reduced State
Figure 7: Block diagram of DSP flows with reduced-state MLSE. The pre-decision is from FFE and the reduced states are used for BM calculation.
To reduce the number of adders and comparators, the reduced state is combined with computational simplification. In our paper, we refer to L2S-MLSE with the combination of computational simplification and reduced state as simplified L2S-MLSE. As shown in Fig. 7, the output signal $d$ of FFE can be regarded as the pre-decision value of the received signal [hoon]. For PAM4 with 4 states of $\{-3,-1,1,3\}$, if the number of states is reduced from 4 to 3 through pre-decision value, the possible states include $\{-3,-1,1\}$ or $\{-1,1,3\}$, which are represented by $\{a,b,c\}$ in Fig. 8(a). If the number of states is reduced from 4 to 2 through pre-decision value, the possible states include $\{-3,-1\}$, $\{-1,1\}$ or $\{1,3\}$, which are represented by $\{a,b\}$ in Fig. 8(b). As shown in Fig. 8(a), when the state number is reduced to 3, the number of ACS units in a 2-step calculation is reduced from 16 to 9, in which only 27 adders and 18 comparators are needed. As shown in Fig. 8(b), when the state number is reduced to 2, the number of ACS units in a 2-step calculation is reduced from 16 to 4, in which only 8 adders and 4 comparators are needed.
1) BM Calculation and ACS of the First Layer: For the first layer, although the reduced state is applied, all the possible state sets can be encountered. Therefore, all the equations in $\bf{A}$}, $\bf{B}$}, $\bf{C}$} and $\bf{D}$} should be calculated in advance. There is no difference for the calculation of $\bf{A}$}, $\bf{B}$}, $\bf{C}$} and $\bf{D}$} no matter whether the reduced state is applied. When calculating ${\bf{A}}+{\bf{B}}$, ${\bf{A}}-{\bf{B}}$, ${\bf{C}}+{\bf{D}}$ or ${\bf{C}}-{\bf{D}}$, the number of adders is reduced to 27 for 3 states and 8 for 2 states. After the adders, the number of comparators is reduced to 18 for 3 states and 4 for 2 states. In conclusion, for 3 states, the computational complexity of the first layer includes $N+1$ variable multipliers, $3N/2+4$ constant multipliers, $34N+31$ adders, and $9N$ comparators. For 2 states, the computational complexity of the first layer includes $N+1$ variable multipliers, $3N/2+4$ constant multipliers, $49N/2+31$ adders, and $2N$ comparators.
Figure 8: Illustration of 2-step trellis in the first layer of L2S-MLSE when the state number is reduced from (a) 4 to 3 or (b) 4 to 2.
2) ACS of the Other Layers and Survivor Path Selection: From the second layer to the last layer, due to the decrease in ACS units, 27 adders and 18 comparators are needed for a 2-step calculation with 3 states. 8 adders and 4 comparators are needed for a 2-step calculation with 2 states. In the final survivor path selection step, the path with the minimal accumulated metrics is chosen from 9 paths for the state number of 3, reducing the number of comparators to 8. Similarly, the path with the minimal accumulated metrics is chosen from 4 paths for the state number of 2, reducing the number of comparators to 3.
Figure 9 compares the BER performance of L2S-MLSE under three configurations: all 4 states (with only computational simplification), simplified L2S-MLSE with 3 states, and simplified L2S-MLSE with 2 states. In our experiment, the noise enhancement caused by FFE is not severe enough to damage the pre-decision value from FFE for 3 or 2 reduced states. Therefore, the BER performance remains nearly identical regardless of whether all 4 states, 3 states, or 2 states are used. For the system with a larger $\alpha$ value, due to the larger noise in the pre-decision value, the BER performance of 2 states may have worse BER performance than 3 states [hoon]. Owing to the nearly identical BER performance in Fig. 9, the simplified scheme in our paper refers to the combination of computational simplification and 2 reduced states. To make a fair comparison, the simplified scheme is also applied to 1S-MLSE. The overall computational complexity of 1S-MLSE, simplified 1S-MLSE, L2S-MLSE, and simplified L2S-MLSE is summarized in Table I.
TABLE I: Computational complexity of 1S-MLSE, simplified 1S-MLSE, L2S-MLSE, and simplified L2S-MLSE
1S-MLSE Simplified 1S-MLSE L2S-MLSE Simplified L2S-MLSE
Variable Multiplier $16N$ $N+1$ $16N$ $N+1$
Constant Multiplier 11 $2N+3$ 11 $3N/2+4$
Adder $32N$ $16N+8$ 80N4880N-48 $57N/2+23$
Comparator $12N+3$ $2N+1$ 48N3348N-33 4N14N-1
Figure 9: BER performance of L2S-MLSE with computational simplification and different numbers of reduced states.
V Experimental Setup
The experiment of a 112-Gbit/s PAM4 signal over 2-km SSMF transmission in C-band is performed to verify the performance of the proposed MLSE. The experimental setup is shown in Fig. 10. At the transmitter, $4\times 10^{5}$ pseudo-random binary sequences (PRBS) are mapped to $2\times 10^{5}$ PAM4 symbols, in which the first 1000 symbols are served as training symbols. To improve the transmission performance, the optimized pre-equalization method proposed in our previous work is applied to compensate for partial bandwidth-limited distortion [guo]. The PAM4 signal is pulse-shaped by a root-raised cosine (RRC) filter with the roll-off factor of 0.1. Then, the signal is resampled by 1.25 sps to align with the sampling rate of the digital-to-analog converter (DAC). The 56-GBaud PAM4 analog signal is generated by a DAC with a sampling rate of 70 GSa/s. The 3-dB bandwidth of the 8-bit DAC is 16 GHz. Afterwards, the analog signal is amplified by a 40-Gbps electrical amplifier (EA, CENTELLAX OA4MVM3). A 40-Gbps optical electro-absorption modulator (EAM, OM5757C-CTM388) is employed to modulate the electrical signal on an optical carrier at $\sim$1550 nm. The DC bias applied to the EAM is -1.6 V, and the insertion loss of the EAM is $\sim$9 dB.
After 2-km SSMF transmission, the received optical power (ROP) is adjusted by a variable optical attenuator (VOA). Then, the optical signal is converted to an electrical signal through a 30-Gbps PIN photodiode with an integrated trans-impedance amplifier (PIN-TIA). Finally, the electrical signal is converted to a digital signal by a 70 GSa/s analog-to-digital converter (ADC). The offline digital signal processing includes matched filter, resampling, synchronization, FFE, post-filter, MLSE, PAM4 demapping, and BER calculation. The estimated frequency domain channel response of this bandwidth-limited system is illustrated in the inset of Fig. 10. Under 2-km SSMF transmission, the 10-dB bandwidth of the whole system is only $\sim$15 GHz. Meanwhile, the frequency response drops rapidly beyond 21 GHz, which leads to severe bandwidth-limited distortion on the PAM4 signal.
Figure 10: Experiment setup of 112-Gbit/s PAM4 based on MLSE after 2-km SSMF transmission.
VI Experimental Results and Discussions
VI-A Parameter Configuration for MLSE
The inputs of MLSE include the coefficients of the post-filter and the signal after the post-filter. For the 2-tap post-filter with the coefficients of $1,\alpha]$, the optimal value of $\alpha$ is swept to obtain the best BER performance after MLSE. The conventional 1S-MLSE is applied to determine the value of $\alpha$, because the conventional 1S-MLSE serves as a performance benchmark for the other proposed MLSE algorithms. After 2-km SSMF transmission, to achieve the best BER performance, the optimal $\alpha$ value is set to 0.55. This value is not too large owing to the use of pre-equalization.
After determining the value of $\alpha$, the number of received symbols $R$ and overlapped symbols $O$ should be determined for each sliding block. For serial MLSE, the PM of the previous block is transferred to the next block. The optimal performance is achieved through the transfer of PM, so that each block only includes received symbols of length $R$ and post-overlap of length $O$. For the parallel MLSE used in our paper, the PM is not transferred to the following blocks. Each block includes pre-overlap of length $O$, received symbols of length $R$, and post-overlap of length $O$. Because the parallel MLSE uses the accumulated BM of pre-overlap to serve as the PM in serial MLSE, the principle to determine the value of $R$ and $O$ is that the parallel MLSE has almost the same BER performance as the conventional serial MLSE. The 1S-MLSE is still applied to determine the value of $R$ and $O$.
Figure 11 illustrates different values of overlapped symbols $O$ for parallel MLSE. In our experiment, the number of received symbols $R$ is set to 16, since smaller values impair the performance of parallel MLSE. When the number of overlapped symbols is equal to or larger than 8, parallel MLSE can achieve almost the same BER performance as serial MLSE. When the number of overlapped symbols is less than 8, parallel MLSE has worse BER performance than serial MLSE. Therefore, in the parallel MLSE architecture of our paper, the number of symbols $N$ in a sliding block is set to 32, which includes 8 symbols of pre-overlap, 16 symbols of useful data, and 8 symbols of post-overlap. All the algorithms described in our paper are applied with this parallel architecture.
Figure 11: BER performance of parallel MLSE with different values of overlapped symbols $O$.
Figure 12: Experimental results of 112-Gbit/s PAM4 with FFE only and simplified L2S-MLSE after 2-km SSMF transmission.
VI-B BER Performance, Latency and Complexity
After determining the parallel architecture, the experimental results of 112-Gbit/s PAM4 after 2-km SSMF transmission are shown in Fig. 12. As described in Section IV-B, the term simplified L2S-MLSE in our paper refers to L2S-MLSE with the combination of computational simplification and 2 reduced states. At the ROP of -7.5 dBm, the BER performance of 112-Gbit/s PAM4 after 2-km SSMF transmission achieves the 7% FEC limit with simplified L2S-MLSE, whereas it fails to meet the 7% FEC limit when only FFE is applied. If the ROP is further increased, the BER performance becomes worse due to the saturation of the TIA. With the symbol length $N=32$, the latency and complexity of 1S-MLSE, simplified 1S-MLSE, L2S-MLSE, and simplified L2S-MLSE are shown in Fig. 13. L2S-MLSE architecture can reduce the latency from 34 delay units in linear order to 7 delay units in logarithmic order.
When considering computational complexity, both constant multipliers and comparators can be implemented using adders. Thus, variable multipliers occupy far more hardware resources than constant multipliers, adders, or comparators. In simplified 1S-MLSE and simplified L2S-MLSE, the number of variable multipliers is reduced from 512 in exponential order to 33 in linear order. Compared with L2S-MLSE, the number of adders and comparators for simplified L2S-MLSE is reduced to 37.2% and 8.4%, respectively. The number of adders and comparators in L2S-MLSE is 2.45 times and 3.88 times that of 1S-MLSE, respectively, while the number of adders and comparators in simplified L2S-MLSE drops to 1.80 times and 1.95 times that of simplified 1S-MLSE, respectively. In conclusion, the simplified scheme can reduce the utilization of multipliers, adders, and comparators.
Figure 13: Latency and computational complexity of 1S-MLSE, simplified 1S-MLSE, L2S-MLSE, and simplified L2S-MLSE with $N=32$.
$BM_{n+1}(-3,-3)=\min\{A_{1}+B_{1},\hskip 7.0ptA_{2}+B_{2},\hskip 7.0ptA_{3}+B_{3},\hskip 7.0ptA_{4}+B_{4}\hskip 8.0pt\}$ (A.1)
$BM_{n+1}(-1,-3)=\min\{A_{5}+B_{5},\hskip 7.0ptA_{6}+B_{6},\hskip 7.0ptA_{7}+B_{7},\hskip 7.0ptA_{8}+B_{8}\hskip 8.0pt\}$ (A.2)
$BM_{n+1}(1,-3)\hskip 7.0pt=\min\{A_{9}+B_{9},\hskip 7.0ptA_{10}+B_{10},A_{11}+B_{11},A_{12}+B_{12}\hskip 1.0pt\}$ (A.3)
$BM_{n+1}(3,-3)\hskip 7.0pt=\min\{A_{13}+B_{13},A_{14}+B_{14},A_{15}+B_{15},A_{16}+B_{16}\hskip 1.0pt\}$ (A.4)
$BM_{n+1}(-3,-1)=\min\{C_{1}+D_{1},\hskip 7.0ptC_{2}+D_{2},\hskip 7.0ptC_{3}+D_{3},\hskip 7.0ptC_{4}+D_{4}\hskip 8.0pt\}$ (A.5)
$BM_{n+1}(-1,-1)=\min\{C_{5}+D_{5},\hskip 7.0ptC_{6}+D_{6},\hskip 7.0ptC_{7}+D_{7},\hskip 7.0ptC_{8}+D_{8}\hskip 8.0pt\}$ (A.6)
$BM_{n+1}(1,-1)\hskip 7.0pt=\min\{C_{9}+D_{9},\hskip 7.0ptC_{10}+D_{10},C_{11}+D_{11},C_{12}+D_{12}\}$ (A.7)
$BM_{n+1}(3,-1)\hskip 7.0pt=\min\{C_{13}+D_{13},C_{14}+D_{14},C_{15}+D_{15},C_{16}+D_{16}\}$ (A.8)
$BM_{n+1}(-3,1)\hskip 7.0pt=\min\{C_{16}-D_{16},C_{15}-D_{15},C_{14}-D_{14},C_{13}-D_{13}\}$ (A.9)
$BM_{n+1}(-1,1)\hskip 7.0pt=\min\{C_{12}-D_{12},C_{11}-D_{11},C_{10}-D_{10},C_{9}-D_{9}\hskip 7.0pt\}$ (A.10)
$BM_{n+1}(1,1)\hskip 14.0pt=\min\{C_{8}-D_{8},\hskip 8.0ptC_{7}-D_{7},\hskip 7.0ptC_{6}-D_{6},\hskip 7.0ptC_{5}-D_{5}\hskip 7.0pt\}$ (A.11)
$BM_{n+1}(3,1)\hskip 14.0pt=\min\{C_{4}-D_{4},\hskip 8.0ptC_{3}-D_{3},\hskip 7.0ptC_{2}-D_{2},\hskip 7.0ptC_{1}-D_{1}\hskip 7.0pt\}$ (A.12)
$BM_{n+1}(-3,3)\hskip 7.0pt=\min\{A_{16}-B_{16},A_{15}-B_{15},A_{14}-B_{14},A_{13}-B_{13}\hskip 1.0pt\}$ (A.13)
$BM_{n+1}(-1,3)\hskip 7.0pt=\min\{A_{12}-B_{12},A_{11}-B_{11},A_{10}-B_{10},A_{9}-B_{9}\hskip 8.0pt\}$ (A.14)
$BM_{n+1}(1,3)\hskip 14.0pt=\min\{A_{8}-B_{8},\hskip 8.0ptA_{7}-B_{7},\hskip 7.0ptA_{6}-B_{6},\hskip 7.0ptA_{5}-B_{5}\hskip 8.0pt\}$ (A.15)
$BM_{n+1}(3,3)\hskip 14.0pt=\min\{A_{4}-B_{4},\hskip 8.0ptA_{3}-B_{3},\hskip 7.0ptA_{2}-B_{2},\hskip 7.0ptA_{1}-B_{1}\hskip 8.0pt\}$ (A.16)
VII Conclusion
In this paper, we propose and experimentally verify the simplified L2S-MLSE algorithm. The simplified scheme combines computational simplification and reduced state. Computational simplification does not degrade the BER performance. The noise enhancement caused by FFE is not severe enough to damage the pre-decision value in our experiment. Therefore, the use of reduced state maintains nearly identical performance. In the experiment of 112-Gbit/s PAM4 transmission under 2-km SSMF, the simplified L2S-MLSE can achieve better BER performance than FFE, enabling the BER performance to reach the 7% FEC limit. Compared with simplified 1S-MLSE, the latency of simplified L2S-MLSE is decreased from 34 delay units in linear order to 7 delay units in logarithmic order. Compared with L2S-MLSE, the simplified L2S-MLSE reduces the number of variable multipliers from 512 in exponential order to 33 in linear order, while the number of adders and comparators is reduced to 37.2% and 8.4%, respectively. In conclusion, the proposed simplified L2S-MLSE shows great potential for future high-speed, low-latency, and low-complexity optical interconnects.
Appendix A
The 64 expansion equations of BM calculation and addition operations for the first layer of L2S-MLSE are shown from Eq. (A.1) to Eq. (A.16). The common term $y_{n}^{2}$ and $y_{n+1}^{2}$ of all the equations are deleted to reduce the computational complexity. In each equation from Eq. (A.1) to Eq. (A.16), among the 4 candidate paths corresponding to the 4 equations, the path with the minimal accumulated BM is selected. The vectors ${\bf{A}}=[{A_{1}},{A_{2}},...,{A_{16}}]$ and ${\bf{C}}=[{C_{1}},{C_{2}},...,{C_{16}}]$ are used to represent the elements related to $\alpha$ and constant values. The vectors ${\bf{B}}=[{B_{1}},{B_{2}},...,{B_{16}}]$ and ${\bf{D}}=[{D_{1}},{D_{2}},...,{D_{16}}]$ are used to represent the other elements, which include the elements related to the received symbol $y$. The detailed definitions of $\bf{A}$}, $\bf{B}$}, $\bf{C}$} and $\bf{D}$} are shown in Table A.I.
Equation (A.1) shows the expansion equations of $BM_{n+1}(-3,-3)$ from 3-3 at time index n1n-1 to 3-3 at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.1). $A_{1}$, $A_{2}$, $A_{3}$, and $A_{4}$ represent the elements related to $\alpha$ and constant values. $B_{1}$, $B_{2}$, $B_{3}$, and $B_{4}$ represent the other elements related to the received symbol $y$.
Equation (A.2) shows the expansion equations of $BM_{n+1}(-1,-3)$ from 1-1 at time index n1n-1 to 3-3 at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.2). $A_{5}$, $A_{6}$, $A_{7}$, and $A_{8}$ represent the elements related to $\alpha$ and constant values. $B_{5}$, $B_{6}$, $B_{7}$, and $B_{8}$ represent the other elements related to the received symbol $y$.
Equation (A.3) shows the expansion equations of $BM_{n+1}(1,-3)$ from $1$ at time index n1n-1 to 3-3 at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.3). $A_{9}$, $A_{10}$, $A_{11}$, and $A_{12}$ represent the elements related to $\alpha$ and constant values. $B_{9}$, $B_{10}$, $B_{11}$, and $B_{12}$ represent the other elements related to the received symbol $y$.
Equation (A.4) shows the expansion equations of $BM_{n+1}(3,-3)$ from $3$ at time index n1n-1 to 3-3 at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.4). $A_{13}$, $A_{14}$, $A_{15}$, and $A_{16}$ represent the elements related to $\alpha$ and constant values. $B_{13}$, $B_{14}$, $B_{15}$, and $B_{16}$ represent the other elements related to the received symbol $y$.
Equation (A.5) shows the expansion equations of $BM_{n+1}(-3,-1)$ from 3-3 at time index n1n-1 to 1-1 at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.5). $C_{1}$, $C_{2}$, $C_{3}$, and $C_{4}$ represent the elements related to $\alpha$ and constant values. $D_{1}$, $D_{2}$, $D_{3}$, and $D_{4}$ represent the other elements related to the received symbol $y$.
Equation (A.6) shows the expansion equations of $BM_{n+1}(-1,-1)$ from 1-1 at time index n1n-1 to 1-1 at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.6). $C_{5}$, $C_{6}$, $C_{7}$, and $C_{8}$ represent the elements related to $\alpha$ and constant values. $D_{5}$, $D_{6}$, $D_{7}$, and $D_{8}$ represent the other elements related to the received symbol $y$.
TABLE A.I: Equations of the vectors $\bf{A}$, $\bf{B}$, $\bf{C}$, $\bf{D}$ from index $1$ to index 1616
Index $\bf{A}$ $\bf{B}$ $\bf{C}$ $\bf{D}$
1 $18{\alpha^{2}}$ $36\alpha +$18+18 $\hskip 7.0pt6y_{n}+6y_{n+1}+6\alpha y_{n}+6\alpha y_{n+1}$ $18{\alpha^{2}}$ $24\alpha +$10+10 $\hskip 7.0pt6y_{n}+2y_{n+1}+6\alpha y_{n}+6\alpha y_{n+1}$
2 $10{\alpha^{2}}$ $12\alpha +$10+10 $\hskip 7.0pt2y_{n}+6y_{n+1}+6\alpha y_{n}+2\alpha y_{n+1}$ $10{\alpha^{2}}$ $8\alpha +$2+2 $\hskip 7.0pt2y_{n}+2y_{n+1}+6\alpha y_{n}+2\alpha y_{n+1}$
3 $10{\alpha^{2}}$ $12\alpha$ +10+10 $2y_{n}+6y_{n+1}+6\alpha y_{n}-2\alpha y_{n+1}$ $10{\alpha^{2}}$ $8\alpha$ +2+2 $2y_{n}+2y_{n+1}+6\alpha y_{n}-2\alpha y_{n+1}$
4 $18{\alpha^{2}}$ $36\alpha$ +18+18 $6y_{n}+6y_{n+1}+6\alpha y_{n}-6\alpha y_{n+1}$ $18{\alpha^{2}}$ $24\alpha$ +10+10 $6y_{n}+2y_{n+1}+6\alpha y_{n}-6\alpha y_{n+1}$
5 $10{\alpha^{2}}$ $24\alpha +$18+18 $\hskip 7.0pt6y_{n}+6y_{n+1}+2\alpha y_{n}+6\alpha y_{n+1}$ $10{\alpha^{2}}$ $12\alpha +$10+10 $\hskip 7.0pt6y_{n}+2y_{n+1}+2\alpha y_{n}+6\alpha y_{n+1}$
6 $2{\alpha^{2}}$ $8\alpha +$10+10 $\hskip 7.0pt2y_{n}+6y_{n+1}+2\alpha y_{n}+2\alpha y_{n+1}$ $2{\alpha^{2}}$ $4\alpha +$2+2 $\hskip 7.0pt2y_{n}+2y_{n+1}+2\alpha y_{n}+2\alpha y_{n+1}$
7 $2{\alpha^{2}}$ $8\alpha$ +10+10 $2y_{n}+6y_{n+1}+2\alpha y_{n}-2\alpha y_{n+1}$ $2{\alpha^{2}}$ $4\alpha$ +2+2 $2y_{n}+2y_{n+1}+2\alpha y_{n}-2\alpha y_{n+1}$
8 $10{\alpha^{2}}$ $24\alpha$ +18+18 $6y_{n}+6y_{n+1}+2\alpha y_{n}-6\alpha y_{n+1}$ $10{\alpha^{2}}$ $12\alpha$ +10+10 $6y_{n}+2y_{n+1}+2\alpha y_{n}-6\alpha y_{n+1}$
9 $10{\alpha^{2}}$ $12\alpha +$18+18 $\hskip 7.0pt6y_{n}+6y_{n+1}-2\alpha y_{n}+6\alpha y_{n+1}$ $10{\alpha^{2}}$ +0+0 +10+10 $\hskip 7.0pt6y_{n}+2y_{n+1}-2\alpha y_{n}+6\alpha y_{n+1}$
10 $2{\alpha^{2}}$ $4\alpha +$10+10 $\hskip 7.0pt2y_{n}+6y_{n+1}-2\alpha y_{n}+2\alpha y_{n+1}$ $2{\alpha^{2}}$ +0+0 +2+2 $\hskip 7.0pt2y_{n}+2y_{n+1}-2\alpha y_{n}+2\alpha y_{n+1}$
11 $2{\alpha^{2}}$ $4\alpha$ +10+10 $2y_{n}+6y_{n+1}-2\alpha y_{n}-2\alpha y_{n+1}$ $2{\alpha^{2}}$ +0+0 +2+2 $2y_{n}+2y_{n+1}-2\alpha y_{n}-2\alpha y_{n+1}$
12 $10{\alpha^{2}}$ $12\alpha$ +18+18 $6y_{n}+6y_{n+1}-2\alpha y_{n}-6\alpha y_{n+1}$ $10{\alpha^{2}}$ +0+0 +10+10 $6y_{n}+2y_{n+1}-2\alpha y_{n}-6\alpha y_{n+1}$
13 $18{\alpha^{2}}$ +0+0 +18+18 $\hskip 7.0pt6y_{n}+6y_{n+1}-6\alpha y_{n}+6\alpha y_{n+1}$ $18{\alpha^{2}}$ $12\alpha$ +10+10 $\hskip 7.0pt6y_{n}+2y_{n+1}-6\alpha y_{n}+6\alpha y_{n+1}$
14 $10{\alpha^{2}}$ +0+0 +10+10 $\hskip 7.0pt2y_{n}+6y_{n+1}-6\alpha y_{n}+2\alpha y_{n+1}$ $10{\alpha^{2}}$ $4\alpha$ +2+2 $\hskip 7.0pt2y_{n}+2y_{n+1}-6\alpha y_{n}+2\alpha y_{n+1}$
15 $10{\alpha^{2}}$ +0+0 +10+10 $2y_{n}+6y_{n+1}-6\alpha y_{n}-2\alpha y_{n+1}$ $10{\alpha^{2}}$ $4\alpha +$2+2 $2y_{n}+2y_{n+1}-6\alpha y_{n}-2\alpha y_{n+1}$
16 $18{\alpha^{2}}$ +0+0 +18+18 $6y_{n}+6y_{n+1}-6\alpha y_{n}-6\alpha y_{n+1}$ $18{\alpha^{2}}$ $12\alpha +$10+10 $6y_{n}+2y_{n+1}-6\alpha y_{n}-6\alpha y_{n+1}$
Equation (A.7) shows the expansion equations of $BM_{n+1}(1,-1)$ from $1$ at time index n1n-1 to 1-1 at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.7). $C_{9}$, $C_{10}$, $C_{11}$, and $C_{12}$ represent the elements related to $\alpha$ and constant values. $D_{9}$, $D_{10}$, $D_{11}$, and $D_{12}$ represent the other elements related to the received symbol $y$.
Equation (A.8) shows the expansion equations of $BM_{n+1}(3,-1)$ from $3$ at time index n1n-1 to 1-1 at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.8). $C_{13}$, $C_{14}$, $C_{15}$, and $C_{16}$ represent the elements related to $\alpha$ and constant values. $D_{13}$, $D_{14}$, $D_{15}$, and $D_{16}$ represent the other elements related to the received symbol $y$.
Equation (A.9) shows the expansion equations of $BM_{n+1}(-3,1)$ from 3-3 at time index n1n-1 to $1$ at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.9). $C_{16}$, $C_{15}$, $C_{14}$, and $C_{13}$ represent the elements related to $\alpha$ and constant values. $D_{16}$, $D_{15}$, $D_{14}$, and $D_{13}$ represent the other elements related to the received symbol $y$.
Equation (A.10) shows the expansion equations of $BM_{n+1}(-1,1)$ from 1-1 at time index n1n-1 to $1$ at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.10). $C_{12}$, $C_{11}$, $C_{10}$, and $C_{9}$ represent the elements related to $\alpha$ and constant values. $D_{12}$, $D_{11}$, $D_{10}$, and $D_{9}$ represent the other elements related to the received symbol $y$.
Equation (A.11) shows the expansion equations of $BM_{n+1}(1,1)$ from $1$ at time index n1n-1 to $1$ at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.11). $C_{8}$, $C_{7}$, $C_{6}$, and $C_{5}$ represent the elements related to $\alpha$ and constant values. $D_{8}$, $D_{7}$, $D_{6}$, and $D_{5}$ represent the other elements related to the received symbol $y$.
Equation (A.12) shows the expansion equations of $BM_{n+1}(3,1)$ from $3$ at time index n1n-1 to $1$ at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.12). $C_{4}$, $C_{3}$, $C_{2}$, and $C_{1}$ represent the elements related to $\alpha$ and constant values. $D_{4}$, $D_{3}$, $D_{2}$, and $D_{1}$ represent the other elements related to the received symbol $y$.
Equation (A.13) shows the expansion equations of $BM_{n+1}(-3,3)$ from 3-3 at time index n1n-1 to 33 at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.13). $A_{16}$, $A_{15}$, $A_{14}$, and $A_{13}$ represent the elements related to $\alpha$ and constant values. $B_{16}$, $B_{15}$, $B_{14}$, and $B_{13}$ represent the other elements related to the received symbol $y$.
Equation (A.14) shows the expansion equations of $BM_{n+1}(-1,3)$ from 1-1 at time index n1n-1 to 33 at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.14). $A_{12}$, $A_{11}$, $A_{10}$, and $A_{9}$ represent the elements related to $\alpha$ and constant values. $B_{12}$, $B_{11}$, $B_{10}$, and $B_{9}$ represent the other elements related to the received symbol $y$.
Equation (A.15) shows the expansion equations of $BM_{n+1}(1,3)$ from $1$ at time index n1n-1 to 33 at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.15). $A_{8}$, $A_{7}$, $A_{6}$, and $A_{5}$ represent the elements related to $\alpha$ and constant values. $B_{8}$, $B_{7}$, $B_{6}$, and $B_{5}$ represent the other elements related to the received symbol $y$.
Equation (A.16) shows the expansion equations of $BM_{n+1}(3,3)$ from $3$ at time index n1n-1 to 33 at time index $n+1$. The state at time index $n$ is 3-3, 1-1, 11 or $3$, corresponding to the four equations in Eq. (A.16). $A_{4}$, $A_{3}$, $A_{2}$, and $A_{1}$ represent the elements related to $\alpha$ and constant values. $B_{4}$, $B_{3}$, $B_{2}$, and $B_{1}$ represent the other elements related to the received symbol $y$.
TABLE B.I: Equations of the vectors $\mathbf{E}$ and $\mathbf{F}$ from index $1$ to index 88
Index $\mathbf{E}$ $\mathbf{F}$
1 $9\alpha^{2}$ $18\alpha +$9+9 $6y_{n}+6\alpha y_{n}$
2 $\alpha^{2}$ $6\alpha +$9+9 $6y_{n}+2\alpha y_{n}$
3 $\alpha^{2}$ $6\alpha$ +9+9 $6y_{n}-2\alpha y_{n}$
4 $9\alpha^{2}$ $18\alpha$ +9+9 $6y_{n}-6\alpha y_{n}$
5 $9\alpha^{2}$ $6\alpha +$1+1 $2y_{n}+6\alpha y_{n}$
6 $\alpha^{2}$ $2\alpha +$1+1 $2y_{n}+2\alpha y_{n}$
7 $\alpha^{2}$ $2\alpha$ +1+1 $2y_{n}-2\alpha y_{n}$
8 $9\alpha^{2}$ $6\alpha$ +1+1 $2y_{n}-6\alpha y_{n}$
$PM_{n}(-3)=\min\{PM_{n-1}(-3)+E_{1}+F_{1},PM_{n-1}(-1)+E_{2}+F_{2},PM_{n-1}(1)+E_{3}+F_{3},PM_{n-1}(3)+E_{4}+F_{4}\}$ (B.1)
$PM_{n}(-1)=\min\{PM_{n-1}(-3)+E_{5}+F_{5},PM_{n-1}(-1)+E_{6}+F_{6},PM_{n-1}(1)+E_{7}+F_{7},PM_{n-1}(3)+E_{8}+F_{8}\}$ (B.2)
$PM_{n}(1)\hskip 7.0pt=\min\{PM_{n-1}(-3)+E_{8}-F_{8},PM_{n-1}(-1)+E_{7}-F_{7},PM_{n-1}(1)+E_{6}-F_{6},PM_{n-1}(3)+E_{5}-F_{5}\}$ (B.3)
$PM_{n}(3)\hskip 7.0pt=\min\{PM_{n-1}(-3)+E_{4}-F_{4},PM_{n-1}(-1)+E_{3}-F_{3},PM_{n-1}(1)+E_{2}-F_{2},PM_{n-1}(3)+E_{1}-F_{1}\}$ (B.4)
Appendix B
The 16 expansion equations of BM calculation and addition operations for 1S-MLSE are shown from Eq. (B.1) to Eq. (B.4). The common term $y_{n}^{2}$ of all the equations is deleted to reduce the computational complexity. In each equation from Eq. (B.1) to Eq. (B.4), among the 4 candidate paths corresponding to the 4 equations, the path with the minimal PM is selected. The vector ${\bf{E}}=[{E_{1}},{E_{2}},...,{E_{8}}]$ represents the elements related to $\alpha$ and constant values. The vector ${\bf{F}}=[{F_{1}},{F_{2}},...,{F_{8}}]$ represents the elements related to the received symbol $y$. The detailed definitions of $\bf{E}$} and $\bf{F}$} are shown in Table B.I.
Equation (B.1) shows the expansion equations of $PM_{n}(-3)$ from 3-3, 1-1, 11 or $3$ at time index n1n-1 to 3-3 at time index $n$. Four states at time index n1n-1 correspond to four equations in Eq. (B.1). $E_{1}$, $E_{2}$, $E_{3}$, and $E_{4}$ represent the elements related to $\alpha$ and constant values. $F_{1}$, $F_{2}$, $F_{3}$, and $F_{4}$ represent the other elements related to the received symbol $y$.
Equation (B.2) shows the expansion equations of $PM_{n}(-1)$ from 3-3, 1-1, 11 or $3$ at time index n1n-1 to 1-1 at time index $n$. Four states at time index n1n-1 correspond to four equations in Eq. (B.2). $E_{5}$, $E_{6}$, $E_{7}$, and $E_{8}$ represent the elements related to $\alpha$ and constant values. $F_{5}$, $F_{6}$, $F_{7}$, and $F_{8}$ represent the other elements related to the received symbol $y$.
Equation (B.3) shows the expansion equations of $PM_{n}(1)$ from 3-3, 1-1, 11 or $3$ at time index n1n-1 to $1$ at time index $n$. Four states at time index n1n-1 correspond to four equations in Eq. (B.3). $E_{8}$, $E_{7}$, $E_{6}$, and $E_{5}$ represent the elements related to $\alpha$ and constant values. $F_{8}$, $F_{7}$, $F_{6}$, and $F_{5}$ represent the other elements related to the received symbol $y$.
Equation (B.4) shows the expansion equations of $PM_{n}(3)$ from 3-3, 1-1, 11 or $3$ at time index n1n-1 to 33 at time index $n$. Four states at time index n1n-1 correspond to four equations in Eq. (B.4). $E_{4}$, $E_{3}$, $E_{2}$, and $E_{1}$ represent the elements related to $\alpha$ and constant values. $F_{4}$, $F_{3}$, $F_{2}$, and $F_{1}$ represent the other elements related to the received symbol $y$.